OpenStudy (anonymous):

given that t = log 8 x, find each of the following in terms of t find: log 4 2x

terenzreignz (terenzreignz):

\[\large t = \log_8 x\]this?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

You know the change of base formula?

OpenStudy (anonymous):

no, i just started to learn this topic

terenzreignz (terenzreignz):

Okay, until further notice, accept this formula... say you have a log \[\large \log_b M\] And you want to change the base to a new one, say, a. then \[\large \log_bM= \frac{\log_aM}{\log_ab}\]

OpenStudy (anonymous):

okay, i get it

terenzreignz (terenzreignz):

Well, you have \[\huge t = \log_8x \] Change the base to 4 instead, following that formula I gave you.

OpenStudy (anonymous):

okay.... done

terenzreignz (terenzreignz):

What did you get?

OpenStudy (anonymous):

\[\log _{4} x \over \log _{4} 8\] is that it?

terenzreignz (terenzreignz):

Very good. \[\huge t = \frac{\log_4 x}{\log_4 8}\] You'd do well to remember that \(\large \log_48\) is just a constant... which can be multiplied to both sides to get...?

OpenStudy (anonymous):

I don't understand

terenzreignz (terenzreignz):

Multiply \(\large \log_4 8 \) on both sides.

OpenStudy (anonymous):

multiply with which one?

terenzreignz (terenzreignz):

on both sides of this equation \[\huge t = \frac{\log_4 x}{\log_4 8}\]

OpenStudy (anonymous):

i don't know how to multiply it. sorry, but can u tell me. ..

terenzreignz (terenzreignz):

\[\huge t\times \color{red}{\log_48} = \frac{\log_4 x}{\log_4 8}\times\color{red}{\log_4 8}\]

OpenStudy (anonymous):

oh, now I get it!

terenzreignz (terenzreignz):

That is good. Can you continue?

OpenStudy (anonymous):

\[t×\log _{4}8=\log _{4} x\]

terenzreignz (terenzreignz):

Good. And now, you have expressed \(\large \log_4 x\) in terms of t. You can do away with that \(\times\) sign

OpenStudy (anonymous):

\[t \log _{4}8=\log _{4} x\]

OpenStudy (anonymous):

still not understand with this